OpenStudy (spyro517):

The equation of line EF is y = 2x + 1. Write an equation of a line parallel to line EF in slope-intercept form that contains point (0, 2). y = 2x − 4 y = 2x + 2 y = negative 1 over 2x − 4 y = 1 over 2x + 2

OpenStudy (anonymous):

Well the slope has to be the same since its parallel

OpenStudy (anonymous):

so the slope is also 2x

OpenStudy (anonymous):

then plug in (0,2) into the equation y = mx+b

OpenStudy (anonymous):

so y = mx + b ------ y=2x+b to find b

OpenStudy (spyro517):

so it would be d then. can you correct me if I am wrong

OpenStudy (anonymous):

y = mx+b y=2x +b 2=2(0)+b b=2 so it would be y=2x+2

OpenStudy (anonymous):

so its b

OpenStudy (spyro517):

I forgot the 0 counts in this kind of equation. Thank you for the correction. It was very helpful.

OpenStudy (anonymous):

no problem